The own-price elasticity of demand for apples is −1.2. If the price of apples falls by 5 per cent, what will happen to the quantity of apples demanded?


a. It will increase 6 per cent.


b.It will increase 4.2 per cent.


c.It will increase 5 per cent.


d.It will fall 3.8 per cent.


1
Expert's answer
2021-11-14T17:42:55-0500

Solution:

The correct answer is a. It will increase by 6 percent.

 

"Price\; elasticity\; of \;demand (PED) =\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; price}"


"-1.2 =\frac{\%\;change\; in\; quantity\; demanded}{-5\% }"


"\%\;change\; in\; quantity\; demanded = -5\%\times-1.2 = 6\%"


% Change in qty demanded = 6%


Quantity demanded will increase by 6%